Accounts receivable from sales transactions were $51,000 at the beginning of the year and $64,000 at the end of the year. Net income reported on the income statement for the year was $105,000. Exclusive of the effect of other adjustments, the cash flows from operating activities to be reported on the statement of cash flows prepared by the indirect method is

  a. $118,000  
  b. $105,000  
  c. $169,000  
  d. $92,000  
Answer #1
Concepts and reason

Cash flow: Cash flow may be defined as the movement of cash. It consists of two categories such as cash inflow and cash outflow. Cash inflow represents revenue or cash earned. Cash outflow represents cash spent.

Cash flow statement: Cash flow statement is a statement that reports the cash inflow in operating, investing, and financing activities of a company for the said stipulated period. This statement represents the sources and applications or uses of cash. Statement of cash flows provides the amount of cash available to meet the obligations of the corporation (liquidity position

Fundamentals

Operating activities under indirect method: It consists of net income, adjustment for non cash expense and change in operating assets and liabilities. When liability is increased it should be added. When asset is increased it should be deducted in the changes in operating assets and liabilities.

Accounts receivable: Accounts receivable is the amount to be received by the company which has sold goods or rendered any services during the period. It is the asset of the company and thus shown under assets in Balance sheet.

Net income: This is the income of a company which reports all the revenues earned and expenses incurred by the company over a specific period of time. It is the income after adjusting the profit and losses. It is also known as operating income and earnings before interest and taxes.

Sales: Sales is an activity of selling the goods in the market which is sold by buyer and purchased by seller. It is the main source of revenue for the company. It is necessary to have consideration for sales.

Calculate increase in accounts receivable:

Increaseinaccountsreceivable}=OpeningbalanceClosingbalance=$64,000$51,000=$13,000\begin{array}{c}\\\left. \begin{array}{l}\\{\rm{Increase in accounts}}\\\\{\rm{ receivable }}\\\end{array} \right\} = {\rm{Opening balance}} - {\rm{Closing balance}}\\\\{\rm{ = \$ 64,000}} - \$ 51,000\\\\ = \$ 13,000\\\end{array}

Therefore, increase in accounts receivable is $13,000.

Determine the cash flows from operating activities on the statement of cash flows prepared by the indirect method.

Cashflowfromoperatingactivities}=(NetincomeIncreaseinaccountsreceivable)=$105,000$13,000=$92,000\begin{array}{c}\\\left. \begin{array}{l}\\{\rm{Cash flow from }}\\\\{\rm{operating activities}}\\\end{array} \right\}{\rm{ = }}\left( {{\rm{Net income}} - {\rm{Increase in accounts receivable}}} \right)\\\\ = \$ 105,000 - \$ 13,000\\\\ = \$ 92,000\\\end{array}

Therefore, cash flow from operating activities is $92,000.

Ans:

Cash flow from operating activities is $92,000.